# CLEME2.0: Towards More Interpretable Evaluation by Disentangling Edits for Grammatical Error Correction

### v1

@@ 5.1 Ablation Studies
To further compare different metrics from a global perspective, we report the average correlations obtained through the exhaustive enumeration of various parameter configurations. Specifically, all possible parameter combinations are attempted, with a step increment of 0.05. From Table 4, we observe that all the correlations are positive, regardless of the applied correction assumptions, evaluation levels, and weighting techniques. Comparing the unweighted and similarity-based weighted results, we conclude that similarity-based weighting significantly promotes human correlations, even on a global scale. Furthermore, corpus-level metrics tend to attain higher average results than sentence-level metrics; nonetheless, sentence-level metrics with optimal parameters surpass their corpus-level counterparts. This suggests that corpus-level metrics may exhibit enhanced robustness concerning parameter selection.
5.2 Exploration of LLM
The results of LLM-based edit weighting are shown in Table 5. The corpus-level results are quite satisfying and are comparable to those of most metrics such as PT-M2 and CLEME. However, the sentence-level outcomes are less satisfactory. This could be due to the fact that LLM assigns error-editing scores on a scale of 1 to 5, which is notably coarser when contrasted with the 0 to 1 continuous scoring scale. Sentence-level scores depend on the mean of the editing scores within a particular sentence. Consequently, even the slightest bias in the scores assigned by the LLM might lead to significant deviations in the sentence-level outcomes.
Although the LLM has had some success, its performance still falls short when compared to the similarity-based weighting. This might be due to the scoring granularity of the 1 to 5 scale provided by the LLM not being sufficiently fine-tuned. In addition, the score heavily relies on the functionality of the LLM, which proves rather unstable.
5.3 Case Study
Table 6 presents examples of CLEME2.0. In the top group, chunk 2 achieves the highest score (0.056), highlighting its significant impact on the sentence. Although "diagnosed" was correctly amended, the omission of "out" persists, rendering the sentence still incorrect. Chunk 4 represents a hit-correction relating to the singular and plural forms in the source sentence and its low score indicates a minimal impact. Chunks 6 and 8 are types of over-correction. Chunk 6 does not change the original meaning, whereas chunk 8 introduces a more severe error due to an incorrect personal pronoun. In the second group, both chunks 3 and 5 exhibit error-corrections, with chunk 5 scoring higher than chunk 3. Chunk 3 involves issues of tense and singular-plural, while chunk 5 presents a more serious error that completely alters the meaning of the sentence. The weighting scores reflect the superiority of the method. For metrics that do not apply weightings, sorts of edits are uniformly assigned, which does not reflect the actual semantics.
6 Conclusion
This paper proposes CLEME2.0, an interpretable evaluation strategy, which are beneficial to reveal crucial characteristics of GEC systems. To address the limitations of traditional reference-based metrics in capturing semantic nuances, we enhance CLEME2.0 using two innovative edit weighting techniques: similarity-based and LLM-based weighting. Through comprehensive experiments and analyses, we validate the efficacy and reliability of our approach. We believe CLEME2.0 will provide a promising perspective on the task of grammatical error correction.
Limitation
Although CLEME2.0 can be extended to other languages, its effectiveness has not been tested in any language other than English. Furthermore, all reference sets utilized in our experiments are derived from the CoNLL-2014 shared task, which is a second language dataset. To demonstrate the robustness of our approaches, further experiments on evaluation datasets encompassing multiple languages and text domains are required. Finally, we strongly advocate for the construction of new GEC evaluation datasets to advance the development of NLP.
Ethics Statement
Appendix A Details about GEC Meta-Evaluation
A.1 Human Rankings
GJG15 ranking.
Grundkiewicz et al. (2015) propose the first large-scale human judgement dataset towards 12 participating systems of the CoNLL-2014 shared task. In this assessment, 8 native speaker are asked to rank the outputs of all the systems from best to worst. Two system ranking lists are generated using Expected Wins (EW) and TrueSkill (TS) respectively. Since all the involved systems are mostly classical systems such as statistical machine translation approaches Junczys-Dowmunt and Grundkiewicz (2014) and classifier-based approaches Rozovskaya et al. (2014), the dataset may not be a ideal test bed for meta-evaluation in the current time.
SEEDA ranking.
Kobayashi et al. (2024b) identify several limitations of GJG15 ranking dataset, and propose a new human ranking dataset called SEEDA. SEEDA consists of corrections with human ratings along two different granularities: edit-based and sentence-based, covering 12 state-of-the-art systems including large language models (LLMs), and two human corrections with different focuses. Three native English speakers participate in the annotation process. Similar to Grundkiewicz et al. (2015), the overall human rankings are derived from TrueSkill (TS) and Expected Wins (EW) based on pairwise judgments.
| # Sents (Length) | 1,312 (23.0) | 1,312 (23.0) | 1,312 (23.0) | 1,312 (23.0) | 1,312 (23.0) | 1,312 (23.0) |
| # Refs (Length) | 2,624 (22.8) | 13,120 (22.9) | 2,624 (23.2) | 2,624 (22.8) | 2,624 (23.0) | 2,624 (22.2) |
| # Edits (Length) | 5,937 (1.0) | 36,677 (1.0) | 4,500 (1.0) | 8,373 (1.1) | 4,964 (0.9) | 11,033 (1.2) |
| # Unchanged Chunks (Length) | 11,174 (4.8) | 93,496 (2.5) | 8,887 (6.3) | 12,823 (3.8) | 10,748 (5.1) | 14,086 (2.9) |
| # Corrected/Dummy Chunks (Length) | 4,994 (1.3) | 26,948 (2.4) | 3,963 (1.2) | 6,305 (1.7) | 4,221 (1.2) | 6,892 (2.6) |
Appendix B Statistics of Reference Datasets
B.1 Baseline Metrics
ERRANT and SentERRANT Bryant et al. (2017) are the most mainstream GEC metrics, which are based on edit matching. They are able to extract edits more accurately, by utilizing the linguistically enhanced Damerau-Levenshtein algorithm.
CLEME and SentCLEME Ye et al. (2023c) are proposed to provide unbiased scores for multi-reference evaluation. Besides, the authors introduce the correction independence assumption, so CLEME can perform based on either traditional correction dependence or correction independence assumptions.
In addition, for the evaluation of SEEDA, we have additionally added the following evaluation methods in accordance with the evaluation methods reported in Kobayashi et al. (2024b):
SOME Yoshimura et al. (2020b): optimizes human evaluation by fine-tuning BERT separately for criteria such as grammaticality, fluency, and meaning preservation.
Appendix C Detailed Results of Evaluation
We list detailed evaluation results of CLEME2.0 on CoNLL-2014 in Table 8.
Appendix D Experimental Details of LLM-based Edit Weighting
Prompt:
As a grammar correction evaluator, you are required to score the corrected editors for each grammatical error. We will give three lines, the first line is the original sentence given, the second line is the modification made to the editor, and the third line is the output form.
The scoring range is 1-5. The larger the score, the more important the editor’s correction is. Correspondingly, the smaller the score, the less important the editor’s correction is.
1 point indicates that this editor’s modification has almost no impact on the original sentence and is dispensable.
2 points indicates that this editorial change has a slight impact.
3 points indicates that this editor’s changes have a certain impact.
4 points indicates that this editorial change is necessary.
5 points indicates that this editing modification is very necessary and of high importance.
The output format is a score of 1 to 5 points.
Next, I will give you a sentence only with an edit. You need to rate each edit in sequence. The desired output is just a score, without any redundant explanation.
Example Input:
Sentence: Nowadays the technologies were improved a lot compared to the last century.
Edit: were => have
Output (1-5):
Example Output:
5
Note that the output must be a number between 1 and 5. Here is the sample:
Due to the strong semantic understanding capabilities of large language models (LLMs) Yang et al. (2020, 2022b, 2022a); Chai et al. (2024), recent work Sottana et al. (2023) has sparked interest in using LLMs for text evaluation, including the evaluation of grammatical error correction. Inspired by this, we utilize LLMs as weighted scorers to assess the importance of each edit. The template for the LLM is shown in Figure 3. For each edit, the constructed sentence contains only one grammatical error, while the other positions are correct. The second line shows the modification of that edit. The LLM is required to determine the necessity of the modified edit and output a score from 1 to 5. A higher score indicates a greater necessity for the edit modification. We do not inform the LLM of the specific types of edits; instead, we let the larger model evaluate the necessity of the modified edits.
D.1 Hit-Correction Edits
Scenario: The hypothesis and reference sentence are consistent.
Focus: The significance of the transition from the source to the hypothesis sentence.
Scoring: A higher score indicates that the edit from source to reference sentence carries substantial importance. Conversely, a lower score suggests that this transition is less crucial.
D.2 Error-Correction Edits
Scenario: The hypothesis and reference sentence are inconsistent.
Focus: The significance of the transition from the hypothesis to the reference sentence.
Scoring: A high score indicates a critical edit, suggesting significant inaccuracies in the hypothesis sentence. A low score implies that the modification is of minimal importance, indicating the hypothesis sentence is either correct or not substantially incorrect.
D.3 Under-Correction Edits
Scenario: The source and hypothesis sentence remain unchanged.
Focus: The importance of modifications from the source to the reference sentence.
Scoring: A high score implies a critical need for the edit, pointing to a severe under-correction. Conversely, a low score indicates that the edit is of lesser importance, suggesting a mild under-correction.
D.4 Over-Correction Edits
Scenario: The source is equivalent to the reference sentence, leading to two distinct situations:
1.
The reference is not an ideal sentence, and the hypothesis sentence is corrected but deemed overcorrected.
2.
The reference is optimal, necessitating no amendments, yet the hypothesis sentence introduces corrections.
Evaluation:
First Situation: Assess the importance () of the edit from the source to the hypothesis sentence. A higher W1 score indicates that the edit is crucial, suggesting imperfections in the reference sentence. Conversely, a lower score suggests that the edit is of minimal importance, rendering the hypothesis’s correction unnecessary.
Second Situation: Examine the significance () of the edit from the hypothesis to the reference sentence. A higher score indicates that the edit is critical, denoting that the hypothesis’s correction was overly aggressive. A lower score implies the edit was unneeded, making the correction by the hypothesis irrelevant.
Formula: The computation of over-correction score is defined as follow:
This score can be either positive or negative. A higher over-correction score signals a less effective performance by the correction system.
By systematically assessing the necessity and importance of different types of edits, we can better understand and improve the performance of grammatical error correction systems.

### v2

@@ C.3 Extra Results of CLEME with Similarity Weighting
To analyze our method from a global viewpoint, we present the average correlations derived from the exhaustive enumeration of possible parameter configurations. We explore all potential parameter combinations with increments of 0.05. Table 11 shows that all correlations are positive, regardless of the correction assumptions, levels of evaluation, or weighting techniques used. By comparing results from unweighted and similarity-based weighted metrics, we determine that similarity-based weighting substantially enhances human correlation on a global level. Additionally, corpus-level metrics generally achieve higher average values compared to sentence-level metrics. However, sentence-level metrics with optimal parameters can outperform their corpus-level equivalents. This implies that corpus-level metrics might demonstrate greater robustness concerning parameter selection.
@@
C.5 Details Results on CoNLL-2014
Table 12 presents a comprehensive evaluation of CLEME2.0 on CoNLL-2014 across all GEC systems. Our method offers a clear and quantitative examination of detailed features of GEC systems, which other automatic metrics cannot provide. For instance, the CAMB system attains the top hit-correction score of 0.271 for CLEME2.0-dep, which shows that about 27.1% of edits by the system are accurate. The wrong-correction score of 0.194 indicates that 19.4% of edits are correctly placed but incorrect, the under-correction score of 0.534 indicates that 53.4% of grammatical errors are overlooked by the system, and the over-correction score of 0.470 suggests that 47.0% of the edits are unnecessary.
As a result, developers and researchers can pinpoint the aspects of their systems that require enhancement. Furthermore, users can select GEC systems that best meet their requirements. For instance, users might opt for a system with a minimal under-correction score in high-stakes situations, as they expect to detect every possible grammatical mistake even though the system might make some unnecessary edits.